OpenStudy (anonymous):

a six sided die is thrown twice. what is the probability that a 5 appeared exactly once in both rolls?

OpenStudy (anonymous):

2/12 maybe

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

1 dice has 6 side in one throw each side has possiblity of 1/6 when throw twice it is same i think

OpenStudy (anonymous):

1/6th * 1/6th = 1/36th, if you're talking about the chances of getting a specific amount...that's a different story, like 7 is the most probable between two dice since it hosts the highest amount of different rolls: 1 + 6 2 + 5 3 + 4 4 + 3 5 + 2 6 + 1 As opposed to 2 or 12 that only offer one combination: 1 + 1 = 2 6 + 6 = 12

OpenStudy (anonymous):

Oh...If you mean what the chances are it would appear only once in two throws...That would be 1/6th * 5/6th I believe, which would be 5/36th

OpenStudy (anonymous):

First time to get 5 probability is 1/6 and sceond time you can get anything except 5 so probability 5/6 Also you can get first time anything except 5, so probability 5/6 and second time you can get 5, so probability 1/6 1/6*5/6+5/6*1/6=10/36=5/18

OpenStudy (anonymous):

1/36

OpenStudy (anonymous):

no